# Proc2Pddl: Open-Domain Planning Representations from Texts

### v1

@@ 5 Conclusion
We present Proc2Pddl, the first open-domain dataset that juxtaposes natural language and planning domain definition language. Our experiments show that LMs find it challenging to model the precondition and effects of actions, leading to both syntactic and semantic errors. We hope our analysis and dataset helps future progress towards integrating the best of LM and formal planning.
@@ Acknowledgements
This work is supported in part by the DARPA KAIROS Program (contract FA8750-19-2-1004), AFRL (contract FA8750-23-C-0507), the Office of the Director of National Intelligence (ODNI) via the IARPA HIATUS Program (contract 2022-22072200005), the NSF (Award 1928631), and gifts from Roblox and Salesforce. Approved for Public Release, Distribution Unlimited. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, ODNI, IARPA, NSF, AFRL, the U.S. Government, or of Roblox or Salesforce. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ Appendix B Sample Data: $\mathbb{T}$, $\mathbb{DF}$, and $\mathbb{PF}$
Appendix C Choice of Input Text
We consider the following choice of wikiHow texts as $\mathbb{T}$.
($\mathbb{T}$ = all): All steps in a wikiHow article.
($\mathbb{T}$ = rel): In Proc2Pddl, each wikiHow article consists of step paragraphs that may or may not be used in defining the actions in the $\mathbb{DF}$. Hence, a mapping between actions and steps is also annotated. This context includes elevant steps to all actions in a $\mathbb{DF}$.
(e.g., Step 1. Find fresh water… Step 2. Collect food… Step 7. Set up camp…)
($\mathbb{T}$ = map): Each action mapped with steps based on the annotated mapping in Proc2Pddl.
(e.g., clean_water: Step 1. Find fresh water…)
($\mathbb{T}$ = sum): One-line summaries of actions annotated in Proc2Pddl.
(e.g., clean_water; boil water to clean it)
The four prompts are increasingly general. Distinguishing from the requirements, the first text conditions demand accurate information identification and extraction, while the last one clearly defines the action but still requires the model to infer implicit entity states. All prompts request an exact translation.
Intrinsic Extrinsic
Model % action acc. $\mathbb{PF}$ solve exact plan
w/o $\mathbb{T}$ (baseline) 13.7 26.3 3.2
$\mathbb{T}$ =sum 15.9 33.7 4.2
$\mathbb{T}$ =sum, CoT 18.1 35.8 6.3
$\mathbb{T}$ =map 11.8 13.7 2.1
$\mathbb{T}$ =map, CoT 8.9 26.3 1.1
$\mathbb{T}$ =rel 11.6 27.4 0.0
$\mathbb{T}$ =rel, CoT 12.2 21.1 4.2
$\mathbb{T}$ =all 12.1 28.4 0.0
$\mathbb{T}$ =all, CoT 12.1 31.6 0.0
Table 5: Performance of GPT-4 using different portions of text $\mathbb{T}$. Metrics include action-wide accuracy, average edit distance of action definitions, the proportion of $\mathbb{PF}$s that can be solved, and the proportion of generated plans that exactly match the gold plans.
Result analysis   As shown in Table 5, in w/o $\mathbb{T}$ setting, fully relying on its implicit knowledge, the model is already capable of inferring PDDL syntactically and semantically. In w/ $\mathbb{T}$ settings, our model shows an ‘U’ performance in terms of the text length. Using a sentence-long description for each action provided by Proc2Pddl, the model achieves the best performance among all, showing a strong deduction ability with the limited but precise NL input. The all setting ensues, which requires the most extraction rather than inference. In contrast, the middle ones (rel/map) with decreasing signal-to-noise ratio lead to worse results, indicating its shortage of extraction-inference trade-off. The signals contain both the described entity states and step relations, explicitly and implicitly. This shortage may come less from the entity states (e.g., fish, spear in hunt_fish), but more from the relation between actions (e.g., make_spear to hunt_fish) which may be expressed in the sum and all settings.

### v2

@@ 6 Conclusion
We present Proc2Pddl, the first open-domain dataset that juxtaposes natural language and planning domain definition language. Our experiments show that ZPD instructions facilitate LMs’ performance, while still find it challenging to translate the precondition and effects of actions. We hope our instruction design, evaluations and dataset help future progress towards integrating the best of LM and formal planning.
@@ Acknowledgements
This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200005. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ Appendix B Sample Data: $\mathbb{T}$, $\mathbb{DF}$, and $\mathbb{PF}$
Appendix C Few-shot Examples
Example 1:
action: slide_straw_over_skewer
text: Slide a segment of plastic straw over the 2 skewers. Cut a segment of a plastic straw that is equal to the width between the wheel wells on your car. Then, slide it onto 1 of the skewers thats attached to a wheel. Do the same with the other skewer.
pddl_action:
(:action slide_straw_over_skewer
:parameters (?straw1 - straw ?straw2 - straw ?skewer1 - skewer ?skewer2 - skewer ?wheel1 - wheel ?wheel2 - wheel)
:precondition (and
(attached ?skewer1 ?wheel1)
(attached ?skewer2 ?wheel2)
(equal_width ?straw1 ?skewer1)
(equal_width ?straw2 ?skewer2))
:effect (and
(over ?straw1 ?skewer1)
(over ?straw2 ?skewer2)))
entity_state_inference:
- action name: slide_straw_over_skewer
- summarize action: The action involves cutting a segment of a plastic straw and sliding it over two skewers attached to wheels. The straw segments should be equal to the width between the wheel wells on a car.
- what is needed to do the action: A segment of a plastic straw, two skewers attached to wheels.
- what is the result: The straw segments are slid over the skewers, which helps in maintaining the structure of the car and providing stability to the wheels.
- entities:
- entity name: plastic straw segment
- before: Not cut and not on skewers
- after: Cut and slid onto skewers
- entity name: skewer
- before: Without straw segments
- after: With straw segments slid onto them
- entity name: wheel
- before: Attached to skewers, without straw segments on the skewers
- after: Attached to skewers, with straw segments on the skewers
Example 2:
action: mix_sand_with_color
text: Close the bags tightly, then shake and squeeze them to mix the sand. Zip the bags shut tightly first. Next, shake the bags to help distribute the food coloring. You can also gently squeeze and massage the sand through the bags. Keep doing this until the color is even. Add more food coloring if the color is too light for you. Keep in mind that the sand will dry lighter.
pddl_action:
(:action mix_sand_with_color
:parameters (?bag - bag ?sand - sand ?color - food_coloring)
:precondition (and (sealed ?bag) (contains ?bag ?sand) (contains ?bag ?color) (not (evenly_colored ?sand)))
:effect (evenly_colored ?sand))’’’,
entity_state_inference:
- action name: mix_sand_with_color
- summarize action: This action involves mixing colored sand by sealing bags containing sand and food coloring, then shaking, squeezing, and massaging the bags until the sand color is evenly distributed. More food coloring may be added if the color is too light, with the understanding that the sand will dry lighter.
- what is needed to do the action: Sealed bags containing sand and food coloring, ability to shake and squeeze the bags.
- what is the result: The sand inside the bags is evenly colored.
- entities:
- entity name: bags
- before: contains uncolored or unevenly colored sand and food coloring
- after: contains evenly colored sand
- entity name: sand
- before: uncolored or unevenly colored
- after: evenly colored
- entity name: food_coloring
- before: present in the bags
- after: mixed evenly with the sand
Example 3:
action: melt_wax
text: Melt candle wax flakes or cubes in a double boiler. Set a can, jar, or another heat-proof container in a pot, and place your solid candle wax (soy or paraffin are most common) inside. Surround the container with water, filling the pot about halfway up the container, then heat the pot on medium heat to double boil the wax to melt it completely, stirring every minute or so to make sure it melts evenly.
pddl_action:
(:action melt_wax
:parameters (?wax ?container ?pot ?heat_source ?water)
:precondition (and (solid ?wax) (in ?wax ?container) (heatproof ?container) (in ?container ?pot) (in ?water ?pot) (cold ?water))
:effect (and (liquid ?wax) (heated ?water)))
entity_state_inference:
- action name: melt_wax
- summarize action: This action involves melting solid candle wax using a double boiler method. The solid wax is placed in a heat-proof container, which is then placed in a pot filled with water. The pot is heated, and the wax is stirred until it melts completely.
- what is needed to do the action: The action requires solid wax, a heat-proof container, a pot, water, and a heat source.
- what is the result: The solid wax is melted into liquid wax.
- entities:
- entity name: wax
- before: solid
- after: liquid
- entity name: container
- before: empty or containing solid wax
- after: containing liquid wax
- entity name: pot
- before: empty or containing water and container with solid wax
- after: containing water and container with liquid wax
- entity name: water
- before: cold or room temperature
- after: heated
- entity name: heat_source
- before: off
- after: on